OpenStudy (anonymous):

someone help with this.. forgot how to do this: Perpendicular to y = -1/4x+2 contains (15,-8) .... i got y = -1/4x-68... is that right...? if not please explain how to do it... and all the steps

OpenStudy (anonymous):

Perpendicular line has a slope of -1/m of the original line. In this case, m = -1/4. So, the perpendicular line has a slope of 4. The line equation will be: y = 4x + b. Since it passes through 15, -8, you can find b and the line equation.

OpenStudy (anonymous):

Yeah i know that.. look this is how i did it: -8 = 4/1(15)+b -8=60+b -60 -60 __________ -68 = b So for my final answer ... when we put it in y = mx+b form.. would it be y = -1/4x - 68 OR y = 4x - 68 ??